OpenStudy (anonymous):

Given ax³+x²+x+b, find the value of a and b if the remainder when divided by (x-1) and (x+1) are 6 & 2

OpenStudy (anonymous):

P(1) = 6 a(1)³+(1)²+(1)+b = 6 a+b+2 = 6 a+b = 4 -(1) < what's this -1 for ? P(-1) = 2 a(-1)³+(-1)²+(-1)+b = 6 -a + b -2= 6 -a + b = 8 -(2) < what's this -2 for ?

OpenStudy (callisto):

It's not -1 and it's not -2 a+b = 4 -(1) => called a+b = 4 as equation (1) Similarly, -a + b = 8 -(2) => called -a+b = 8 as equation (2)

OpenStudy (anonymous):

Ohhhh ! alright

OpenStudy (anonymous):

do i make (1) = (2) now ?

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

Nope, not really. Solve: a+b = 4 -(1) -a + b = 8 -(2) (1) + (2) What do you get?

OpenStudy (anonymous):

b=-2

OpenStudy (callisto):

How did you get that?

OpenStudy (anonymous):